# FPGA-Enabled Machine Learning Applications in Earth Observation: A Systematic Review

### v1

@@ 4. From AI Models to FPGA Deployment: Design Strategies and Challenges
ML S (Boyle et al., 2023) ZU7EV Deep Belief Network f32 - - 10 0 250 - - - 5.4e-3 -
GNN F (Zhang et al., 2023a) U280 GNN f32* - - 43 37 260 400 38.0 10.5 2.7e-2 759
(Zhang et al., 2022) ZU7EV GNN f32 2.13M 1.0 74 91 125 - 6.3 - 1.1e-4 9500
Automatic FINN CNN S (Hashimoto et al., 2019) 7Z020 CNN b - - - - - - - - 3.3e-1 -
(Myojin et al., 2020) 7Z045 BNN b - - - - - - - - 2.1e-2 -
(Pitonak et al., 2022) 7Z020 CloudSatNet-1 Q2 b - 1.4 14 - 100 - 2.5 - - 15.5
7Z020 CloudSatNet-1 Q4 i4 - 3.1 14 - 100 - 2.6 - - 15.5
Vitis AI CNN F (Huang et al., 2024) ZU19EG ETAUS i8 - - - - 106 329.9 1.6 199.9 1.1e-1 9.3
(Li et al., 2023) ZU7EV RFA-YOLO i8 - 24.8 - - 200 - 15.8 - - 28
(Nguyen et al., 2024) ZU7EV YOLOv4 i8 105 G 245.0 - - 100 - 30.1 - 1.1e-1 17.9
ZU7EV YOLOv4-tiny 3L i8 15.3G 24.1 - - 100 - 26.4 - 1.9e-2 125
(Sabogal and George, 2021) ZU3EG ENet i8 4.06G 0.4 70 75 300 - 3.4 - - 25.2
ZU3EG ESPNet i8 3.71G 0.3 70 75 300 - 3.1 - - 11.7
ZU3EG FPN i8 17.3G 5.8 70 75 300 - 4.0 - - 14.1
ZU3EG U-Net i8 96.6G 7.4 70 75 300 - 3.4 - - 2.7
(Upadhyay et al., 2024) ZU9EG ResNet-50 i8 - - - - - - - - 1.1e-2 93
ZU9EG SICNet i8 - - - - - - - - 1.8e-3 530.7
(Yu et al., 2024) ZU7EV Impr. YOLOv4-tiny i8 - - - - 300 - 20.0 - - 295.9
(Zhao et al., 2023) K26 YOLOv4-MobileNetv3 i8 - 5.7 59 - 200 - 7.2 - - 48.1
@@
ML S (Gyaneshwar and Nidamanuri, 2022) 7A35T CAG-SC2S i32 - -? 93 21 70 - 0.3 - 5.3e-6* -
7A35T CAL-SC2S i32 - -? 96 25 65 - 0.4 - 5.7e-6* -
7A35T SVM i32 - -? 49 6 72 - 0.2 - 5.1e-6* -
VGT CNN S (Bahl et al., 2019) 5CSXC6 C-FCN++ i16 4.70K 0.0 - - 100 - - - 1.5e-1 7
Unclassified - CNN S (Chen et al., 2020b) 7K325T LeNet-5 f32 - - - 33 100 - - - 2.3e-3 -
F (Suh et al., 2021) ZU3EG SSD 0.25x i8 4.04G - 73 47 200 138 2.4 57.5 - 34.2
ZU3EG SSD 0.5x i8 15.6G - 73 55 200 150 2.6 57.7 - 9.6
ZU3EG SSD 1.0x i8 61.6G - - - 200 158 2.0 79.0 - 2.6
- (Wang et al., 2024) 7Z020 Lightweight CNN i8 4.60G - 70 25 100 - 3.3 - - 60
ML S (Hammoud et al., 2022) 7Z020 MLP f32* - - 26 1 - - 0.1 - - -
@@
FPGA: AMD/XILINX FPGA names are stripped of leading XC to increase readability
@@
The designs are split between manual implementations—Hardware Description Language (HDL) and High-Level Synthesis (HLS)—and automatic frameworks—FINN (Umuroglu et al., 2017), Vitis AI (AMD, 2019), MATLAB, and VGT. HDL is the most commonly used language ($\approx 45\%$), offering precise control of the generated circuit. However, most designs do not exploit this level of control to use all available resources nor push the theoretical limit of the clock frequency. High-Level Synthesis (HLS) tools offer a more abstract approach to FPGA programming as they enable the use of high-level languages to create larger designs with less lines of code. Even with HLS, a solid grasp of FPGA architecture and hardware concepts remains necessary, especially when compared to GPU programming environments, like CUDA. In the reported designs, we detect no overarching trend in worse resource utilization or clock frequency of HLS designs compared to HDL designs. The average resource utilization of implementations are $\approx$30DSPs and 50% BRAM for HDL and 46% DSPs and 43% for HLS with a high variance between individual designs. Most manual designs focus on CNN-based networks ($\approx$ 71%), a well-studied network architecture commonly implemented using an array of Processing Elements (PEs) that pass intermediate results between each other.
@@
We split each workload into two distinct design patterns. Flexible (F) designs are bitstreams that can support multiple different kernels and network topologies without requiring reconfiguration. Supporting different networks and kernels usually requires weights to be loaded from off-chip memory. Zhang et al. (2022) use a Flexible design with on-chip memory, pre-loading memory to the weight buffer, and amortizing the access over multiple classifications. The support for different layers comes at a small cost in computational performance, but allows implementing larger models that would otherwise not fit on the FPGA using a Specific approach. The average footprint of the $20$ Flexible designs is $25MB$ compared to $15MB$ for the $10$ Specific designs. Unlike Flexible designs, Specific designs only support a single network architecture. Deploying a different model requires re-implementation of the bitstream or is not supported. This allows the design to be hyper-specialized to the model, making full use of the customizability of the FPGA. Specific designs can use different datapaths for different pruning and quantization of individual layers. Differentiating model families (’Fam.’) is relevant as different ML models have vastly distinct requirements. For example, CNNs spend most of their time performing convolution, a task with high computational intensity (Rapuano et al., 2021), whereas GNNs require a high memory bandwidth (Zhang et al., 2023a). The FPGA device also influences the performance, similar to CPUs, the number of transistors increases over the generations, enabling designers to use more resources such as DSPs, BRAM, and LUTs (see 3.4).
@@
The ’Prec.’ column reports the precision of operations post-quantization. The most popular quantization is i8, with $51\%$ of all implementations and $44\%$ of manual ones using either i8 or a mixture of i8 with f32 or i32. Next, we report the model complexity ’C[OP], an essential metric as a higher complexity requires more computational resources to produce the same temporal throughput ’FPS’. In addition, the relationship between model complexity and memory footprint can indicate whether a model is compute- or memory-bound. Most of the designs do not reach the compute boundary nor provide an off-chip memory bandwidth analysis making a further discussion impossible.
The next metric is the model footprint and the location of parameters. FPGAs, similarly to CPUs, contain restricted on-chip memory, e.g., $<5$MB for Zynq MPSoC chips, and a main off-chip memory ($512$MB to $32$GB). Using on-chip memory eliminates most memory bandwidth concerns as distributed BRAM provides very high bandwidth. On-chip memory weights can either be stored in BRAM—pre-loaded before execution—or stored inside LUTs—disabling any change in parameters. Specific designs mostly use on-chip memory with only $25\%$ using off-chip memory, while $93\%$ of Flexible designs use off-chip memory. In particular, the FINN compiler stores weights on-chip, such as BRAM and LUTs, while Vitis AI loads weights from external memory. Storing the weights on-chip leads to the design being mostly limited by BRAM (Wang and Qiu, 2019; Zhang et al., 2023b; Fraczek et al., 2018). The average memory footprint of on-chip implementation is only $0.83MB$ compared to $28MB$ for off-chip memory implementation.
The last two middle columns present the DSP & BRAM utilization. A high utilization can indicate a good selection of the model for the FPGA, while a low utilization can indicate bad scalability of the design. This excludes network architectures that cannot use DSPs and BRAMs efficiently, such as BNNs—using LUTs for most operations—or GNNs—with low data locality and computational intensity. Yahiaoui et al. (2017) compare an implementation using only LUTs with a DSP- and BRAM-based implementation. They show that the BRAM_DSP implementation can double the number of pixel processed in parallel, while using the same amount of resources.
@@
Comparing different models on CPUs and GPUs is already difficult; for FPGAs, where the design can vary even more, finding a fair comparison is especially complex. One common metric relevant for comparison is the computational throughput, marking the peak number of operations per second. Nevertheless, operations with different precisions cannot be directly compared, as floating-point computations are more expensive than integer operations, and smaller bit widths yield higher [GOP/s]. For example, the highest throughput is reported as $3$ TOP/s (Zhang et al., 2023b). However, the solution uses an All Adder NN, which approximates the functionality of a convolution filter by using only additions. Moreover, the intermediate feature maps resulting from the convolution layers are sent back to the CPU. Thus, the Fully-Connected (FC) layers, corresponding to $92\%$ of the parameters, are executed on the CPU, drastically reducing the memory footprint. The remaining layers only represent $10M$ parameters, which can fit on the boards $6.6$MB BRAM due to the quantization to 4-bit. Using BRAM eliminates any memory bottlenecks and the main computation, a 4-bit adder, can be implemented using solely LUTs, resulting in the extreme reported throughput. The second highest reported throughput is $452$ GOP/s (Li et al., 2019). Compared to Zhang et al. (2023b) the design uses 16-bit values loaded from memory, and the main operations are Multiply-Accumulate (MAC), counted as two operations. The DSP block inside a 7-series Xilinx FPGA can perform one MAC operation per cycle and they manage to keep $1152$ DSPs ($60\%$) busy achieving $1.97\times 1152\times 200MHz\approx 452GOP/s$. Sabogal and George (2021) and Suh et al. (2021) use double datarate DSPs to increase the performance—an optional optimization inside Vitis AI. Doubling the clock frequency of DSP components, which can be clocked higher than the rest of the design, roughly doubles the available computational capability.
A majority of experiments ($\approx 72\%$) report the power consumed by the FPGA, a value pointless on its own as the total energy required to compute a sample is more important. To attempt a more informative comparison between different models and implementations, we calculate the computational efficiency ’T/P’ in $[\frac{GOP/s}{W}]$. Zhang et al. (2023b) achieve the highest computational efficiency $368\frac{GOP/s}{W}$ due to the high throughput detailed previously. On the other hand, He et al. (2023) have an extremely low $<2~{}\frac{GOP/s}{W}$ efficiency, even though they use a low-precision datatype (i8). Another key design metric is the frequency ’MHz’, which depends on the speed-grade and technology of the device, as well as the critical paths in the design. A low frequency, compared to designs on the same FPGA, can hint at a poorly designed circuit with long critical paths. ’Latency’ and ’FPS’ values are model- and implementation-specific, and uncomparable between different workloads and implementations. Nevertheless, we include them in Table 3 to ensure completeness. It is important to separate ’Latency’ from ’FPS’, as, generally, multiple inputs can be processed inside the accelerator pipeline increasing the FPS significantly compared to a non-pipelined implementation that can process one element at a time.
@@
Extensive work in implementing CNNs on FPGAs has produced various automated frameworks supporting common layers. Within the surveyed studies we could not find a trend showing better performance of manual designs in any of the metrics. Specific design patterns are commonly used ($\approx$ 75%) with limited on-chip memory, while Flexible design patterns tend to support larger models (+ 10 MB in average). 2.1 TAKEAWAYS
@@
Instruction Level Parallelism (ILP) exploits the execution of consecutive instructions in parallel, for example overlapping load and MAC operations (Suh et al., 2021). This level of parallelism is exploited in almost every design and can be performed manually in HDL or automatically using HLS control directives for loop statements. Consecutive iterations of a loop can overlap, increasing the throughput and amortizing the latency over multiple loop iterations. In addition, ILP designs commonly exploit Data Level Parallelism (DLP), for example by executing multiple MAC operations in parallel to perform convolution (Yan et al., 2022).
After optimizing individual tasks using ILP, designers can leverage Task Level Parallelism (TLP), for example, by overlapping different layers in a NN—a mechanic heavily exploited by FINN (Umuroglu et al., 2017). The compiler connects individual layers with First-In First-Out (FIFO) streams and starts executing a layer as soon as its first inputs are produced by the previous layer. Increasing the batch size increases TLP, enabling different inputs to be processed at the same time. The disadvantage of larger batch sizes is a possible increase in latency, leading to most implementations using a batch size of $1$ during inference. Increasing computation puts pressure on the memory bandwidth, and most designs use multiple PEs in a Systolic Array (Kung, 1982) to reduce off-chip memory accesses. Furthermore, multiple PEs enable simple scaling to use most of the FPGA resources; this approach can be found in many designs (Chellaswamy et al., 2024; He et al., 2023; Li et al., 2019; Yang et al., 2022; Pitonak et al., 2022; Rapuano et al., 2021; Shibi and Gayathri, 2021; Suh et al., 2021; Wang and Qiu, 2019; Wei et al., 2019; Wu et al., 2021; Yahiaoui et al., 2017; Yan et al., 2022; Yang et al., 2023; Zhang et al., 2022, 2023a, 2023b, 2021, 2020).
@@
Using multiple levels of parallelism, most designs can be scaled to reach the computational limit for floating-point operations on FPGAs, which require multiple DSPs per operation. To increase the OP/s available, designers can choose to reduce datatypes precision, hence reducing the number of resources required per operation. Most designs use quantization to a fixed-point datatype to increase the number of available operations. In addition to DSPs, which can perform basic arithmetic operations on integers, LUTs can perform any 6-bit to 1-bit mapping or a 5-bit to 2-bit operation. Thus, a BNN or low-precision network can be mapped to LUTs directly, allowing for high peak OP/s (Zhang et al., 2023b). The trade-off between LUT and DSP utilization is automatically explored by Yang et al. (2022). In addition to computational benefits, lower precision reduces off-chip memory requirements and can allow the network to fit entirely into BRAM. Another option to increase available operations is to run DSPs at double the frequency, effectively doubling their throughput (Suh et al., 2021; Sabogal and George, 2021). For networks with low computational intensity, the focus is placed on creating an advanced network to reduce the number of access to off-chip memory as much as possible (Zhang et al., 2023a, 2022).

### v2

@@ 4. From AI Models to FPGA Deployment: Design Strategies and Challenges
ML S (Boyle et al., 2023) ZU7EV Deep Belief Network f32 - - 10 0 250 - - - 5.4 ms -
GNN F (Zhang et al., 2023) U280 GNN f32* - - 43 37 260 400 38.00 10.5 27 ms 759
(Zhang et al., 2022) ZU7EV GNN f32 2.13M 0.96 74 91 125 - 6.30 - 0.105 ms 9500
ViT S (Wickramasinghe et al., 2024) U250 VTR - 18.2G -? - - 300 - - - - -
Automatic FINN CNN S (Hashimoto et al., 2019) 7Z020 CNN b - - - - - - - - 330 ms -
(Myojin et al., 2020) 7Z045 BNN b - - - - - - - - 21.14 ms -
(Pitonak et al., 2022) 7Z020 CloudSatNet-1 Q2 b - 1.43 14 - 100 - 2.52 - - 15.46
7Z020 CloudSatNet-1 Q4 i4 - 3.06 14 - 100 - 2.59 - - 15.47
Vitis AI CNN F (Huang et al., 2024) ZU19EG ETAUS i8 - - - - 106 329.9 1.65 199.9 107.8 ms 9.3
(Li et al., 2023a) ZU7EV RFA-YOLO i8 - 24.75 - - 200 - 15.82 - - 27.97
(Nguyen et al., 2024) ZU7EV YOLOv4 i8 105 G 245 - - 100 - 30.1 - 107.94 ms 17.9
ZU7EV YOLOv4-tiny 3L i8 15.3G 24.10 - - 100 - 26.40 - 19.18 ms 125
(Sabogal and George, 2021) ZU3EG ENet i8 4.06G 0.36 70 75 300 - 3.36 - - 25.2
ZU3EG ESPNet i8 3.71G 0.33 70 75 300 - 3.08 - - 11.7
ZU3EG FPN i8 17.3G 5.84 70 75 300 - 4.03 - - 14.1
ZU3EG U-Net i8 96.6G 7.40 70 75 300 - 3.38 - - 2.7
(Upadhyay et al., 2024) ZU9EG ResNet-50 i8 - - - - - - - - 10.7 ms 93.02
ZU9EG SICNet i8 - - - - - - - - 1.8 ms 530.66
(Yu et al., 2024) ZU7EV Improved YOLOv4-tiny i8 - - - - 300 - 20.00 - - 295.9
(Zhao et al., 2023) K26 YOLOv4-MobileNetv3 i8 - 5.69 59 - 200 - 7.20 - - 48.14
@@
ML S (Gyaneshwar and Nidamanuri, 2022) 7A35T CAG-SC2S i32 - -? 93 21 70 - 0.31 - 5.27 us* -
7A35T CAL-SC2S i32 - -? 96 25 65 - 0.36 - 5.70 us* -
7A35T SVM i32 - -? 49 6 72 - 0.22 - 5.12 us* -
VGT CNN S (Bahl et al., 2019) 6 C-FCN++ i16 4.70K 0.047 - - 100 - - - 150 ms 7
Unclassified - CNN S (Chen et al., 2020b) 7K325T LeNet-5 f32 - - - 33 100 - - - 2.29 ms -
F (Suh et al., 2021) ZU3EG SSD 0.25x i8 4.04G - 73 47 200 138 2.40 57.5 - 34.18
ZU3EG SSD 0.5x i8 15.6G - 73 55 200 150 2.60 57.7 - 9.6
ZU3EG SSD 1.0x i8 61.6G - 73 85 200 158 2.00 79.0 - 2.56
- (Wang et al., 2024) 7Z020 Lightweight CNN i8 4.60G - 70 25 100 - 3.30 - - 60.0
ML S (Hammoud et al., 2022) 7Z020 MLP f32* - - 26 1 - - 0.13 - - -
(Hammoud et al., 2024) 7Z020 TinyML i8,i16 - 0.000037 49 - 100 - 0.173 - - -
@@
FPGA: AMD/XILINX Fpga names are stripped of leading to increase readability
@@
The designs are split between manual implementations—Hardware Description Language (HDL) and High-Level Synthesis (HLS)—and automatic frameworks—FINN (Umuroglu et al., 2017), Vitis AI (AMD, 2019), MATLAB, and VGT. HDL is the most commonly used language ($\approx 45\%$), offering precise control of the generated circuit. However, most designs do not exploit this level of control to use all available resources nor push the theoretical limit of the clock frequency. High-Level Synthesis (HLS) tools offer a more abstract approach to FPGA programming as they enable the use of high-level languages to create larger designs with less lines of code. Even with HLS, a solid grasp of FPGA architecture and hardware concepts remains necessary, especially when compared to GPU programming environments, like CUDA. In the reported designs, we detect no overarching trend in worse resource utilization or clock frequency of HLS designs compared to HDL designs. The average resource utilization of implementations are $\approx$30% DSPs and 50% BRAM for HDL and 46% DSPs and 43% for HLS with a high variance between individual designs. Most manual designs focus on CNN-based networks ($\approx$ 70%), a well-studied network architecture commonly implemented using an array of Processing Elements (PEs) that pass intermediate results between each other.
@@
We split each workload into two distinct design patterns. Flexible (F) designs are bitstreams that support multiple different kernels and network topologies without requiring reconfiguration. Supporting different networks and kernels usually requires weights to be loaded from off-chip memory. Zhang et al. (2022) use a Flexible design with on-chip memory, pre-loading memory to the weight buffer, and amortizing the access over multiple classifications. The support for different layers comes at a small cost in computational performance, but allows implementing larger models that would otherwise not fit on the FPGA using a Specific approach. The average footprint of the $20$ Flexible designs is $25MB$ compared to $14MB$ for the $11$ Specific designs. Unlike Flexible designs, Specific designs only support a single network architecture. Deploying a different model requires re-implementation of the bitstream or is not supported. This allows the design to be hyper-specialized to the model, making full use of the customizability of the FPGA. Specific designs can use different datapaths for different pruning and quantization of individual layers. Differentiating model families (’Fam.’) is relevant as different ML models have vastly distinct requirements. For example, CNNs spend most of their time performing convolution, a task with high computational intensity (Rapuano et al., 2021), whereas GNNs require a high memory bandwidth (Zhang et al., 2023). The FPGA device also influences the performance, similar to CPUs, the number of transistors increases over the generations, enabling designers to use more resources such as DSPs, BRAM, and LUTs (see 3.4).
@@
The ’Prec.’ column reports the precision of operations post-quantization. The most popular quantization is i8, with $50\%$ of all implementations and $43\%$ of manual ones using either i8 or a mixture of i8 with f32 or i32. Next, we report the model complexity ’C[OP], an essential metric as a higher complexity requires more computational resources to produce the same temporal throughput ’FPS’. In addition, the relationship between model complexity and memory footprint can indicate whether a model is compute- or memory-bound. Most of the designs do not reach the compute boundary nor provide an off-chip memory bandwidth analysis making a further discussion impossible. The next metric is the model footprint and the location of parameters. FPGAs, similarly to CPUs, contain restricted on-chip memory, e.g., <5<5MB for Zynq MPSoC chips, and a main off-chip memory (512512MB to 3232GB). Using on-chip memory eliminates most memory bandwidth concerns as distributed BRAM provides very high bandwidth. On-chip memory weights can either be stored in BRAM—pre-loaded before execution—or stored inside LUTs—disabling any change in parameters. Specific designs mostly use on-chip memory with only $24\%$ using off-chip memory, while $93\%$ of Flexible designs use off-chip memory. In particular, the FINN compiler stores weights on-chip, such as BRAM and LUTs, while Vitis AI loads weights from external memory. Storing the weights on-chip leads to the design being mostly limited by BRAM (Wang and Qiu, 2019; Zhang et al., 2023; Fraczek et al., 2018). The average memory footprint of on-chip implementation is only $0.72MB$ compared to $28MB$ for off-chip memory implementation. The last two middle columns present the DSP & BRAM utilization. A high utilization can indicate a good selection of the model for the FPGA, while a low utilization can indicate bad scalability of the design. This excludes network architectures that cannot use DSPs and BRAMs efficiently, such as BNNs—using LUTs for most operations—or GNNs—with low data locality and computational intensity. Yahiaoui et al. (2017) compare an implementation using only LUTs with a DSP- and BRAM-based implementation. They show that the BRAM_DSP implementation can double the number of pixel processed in parallel, while using the same amount of resources.
@@
Comparing different models on CPUs and GPUs is already difficult; for FPGAs, where the design can vary even more, finding a fair comparison is especially complex. One common metric relevant for comparison is the computational throughput, marking the peak number of operations per second. Nevertheless, operations with different precisions cannot be directly compared, as floating-point computations are more expensive than integer operations, and smaller bit widths yield higher [GOP/s]. For example, the highest throughput is reported as $3$ TOP/s (Zhang et al., 2023). However, the solution uses an All Adder NN, which approximates the functionality of a convolution filter by using only additions. Moreover, the intermediate feature maps resulting from the convolution layers are sent back to the CPU. Thus, the Fully-Connected (FC) layers, corresponding to $92\%$ of the parameters, are executed on the CPU, drastically reducing the memory footprint. The remaining layers only represent $10M$ parameters, which can fit on the boards 6.66.6MB BRAM due to the quantization to 4-bit. Using BRAM eliminates any memory bottlenecks and the main computation, a 4-bit adder, can be implemented using solely LUTs, resulting in the extreme throughput. The second highest reported throughput is $452$ GOP/s (Li et al., 2019b). Compared to Zhang et al. (2023) the design uses 16-bit values loaded from memory, and the main operations are Multiply-Accumulate (MAC), counted as two operations. The DSP block inside a 7-series Xilinx FPGA can perform one MAC operation per cycle and they manage to keep $1152$ DSPs ($60\%$) busy achieving $1.97\times 1152\times 200MHz\approx 452GOP/s$. Sabogal and George (2021) and Suh et al. (2021) use double datarate DSPs to increase the performance—an optional optimization inside Vitis AI. Doubling the clock frequency of DSP components, which can be clocked higher than the rest of the design, roughly doubles the available computational capability. A majority of experiments ($\approx 72\%$) report the power consumed by the FPGA, a value pointless on its own as the total energy required to compute a sample is more important. To attempt a more informative comparison between different models and implementations, we calculate the computational efficiency ’T/P’ in $\frac{GOP/s}{W}]$. Zhang et al. (2023) achieve the highest computational efficiency $368\frac{GOP/s}{W}$ due to the high throughput detailed previously. On the other hand, He et al. (2023) have an extremely low $2~\frac{GOP/s}{W}$ efficiency, even though they use a low-precision datatype (i8). Another key design metric is the frequency ’MHz’, which depends on the speed-grade and technology of the device, as well as the critical paths in the design. A low frequency, compared to designs on the same FPGA, can hint at a poorly designed circuit with long critical paths. ’Latency’ and ’FPS’ values are model- and implementation-specific, incomparable between different workloads and implementations, and only included to ensure completeness. It is important to separate ’Latency’ from ’FPS’, as, generally, multiple inputs can be processed inside the accelerator pipeline, increasing the FPS significantly compared to a non-pipelined implementation that can process one element at a time.
@@
Extensive work in implementing CNNs on FPGAs has produced various automated frameworks supporting common layers. Within the surveyed studies we could not find a trend showing better performance of manual designs in any of the metrics. Specific design patterns are commonly used ($\sim\!75\%$) with limited on-chip memory, while Flexible design patterns tend to support larger models (+10+10MB in average). 2.1 TAKEAWAYS
@@
Instruction Level Parallelism (ILP) exploits the execution of consecutive instructions in parallel, for example overlapping load and MAC operations (Suh et al., 2021). This level of parallelism is exploited in almost every design and can be performed manually in HDL or automatically using HLS control directives for loop statements. Consecutive iterations of a loop can overlap, increasing the throughput and amortizing the latency over multiple loop iterations. In addition, ILP designs commonly exploit Data Level Parallelism (DLP), for example by executing multiple MAC operations in parallel to perform convolution (Yan et al., 2022). After optimizing individual tasks using ILP, designers can leverage Task Level Parallelism (TLP), for example, by overlapping different layers in a NN—a mechanic heavily exploited by FINN (Umuroglu et al., 2017). The compiler connects individual layers with First-In First-Out (FIFO) streams and starts executing a layer as soon as its first inputs are produced by the previous layer. Increasing the batch size increases TLP, enabling different inputs to be processed at the same time. The disadvantage of larger batch sizes is a possible increase in latency, leading to most implementations using a batch size of $1$ during inference.
@@
Increasing computation puts pressure on the memory bandwidth, and most designs use multiple PEs in a Systolic Array (Kung, 1982) to reduce off-chip memory accesses. Furthermore, transformer-based networks (Wickramasinghe et al., 2024), relying on dense matrix multiplication with a lower computational intensity than convolutional layers, require more off-chip bandwidth. Similarly, GNNs, exhibiting the lowest computational intensity, necessitate an even greater focus on creating an advanced network to reduce the number of accesses to off-chip memory as much as possible (Zhang et al., 2023, 2022). Given enough bandwidth, multiple PEs enable simple scaling to use most of the FPGA resources; this approach can be found in many designs (Chellaswamy et al., 2024; He et al., 2023; Li et al., 2019b; Yang et al., 2022; Pitonak et al., 2022; Rapuano et al., 2021; Shibi and Gayathri, 2021; Suh et al., 2021; Wang and Qiu, 2019; Wei et al., 2019; Wu et al., 2021; Yahiaoui et al., 2017; Yan et al., 2022; Yang et al., 2023; Zhang et al., 2022, 2023, 2023, 2021, 2020). Using multiple levels of parallelism, most designs can be scaled to reach the computational limit for floating-point operations on FPGAs, which require multiple DSPs per operation. To increase the OP/s available, designers can choose to reduce datatypes precision, hence reducing the number of resources required per operation. Most designs use quantization to a fixed-point datatype to increase the number of available operations. In addition to DSPs, which can perform basic arithmetic operations on integers, LUTs can perform any 6-bit to 1-bit mapping or a 5-bit to 2-bit operation. Thus, a BNN or low-precision network can be mapped to LUTs directly, allowing for high peak OP/s (Zhang et al., 2023). The trade-off between LUT and DSP utilization is automatically explored by Yang et al. (2022). In addition to computational benefits, lower precision reduces off-chip memory requirements and can allow the network to fit entirely into BRAM. Another option to increase available operations is to run DSPs at double the frequency, effectively doubling their throughput (Suh et al., 2021; Sabogal and George, 2021).
Finally, the large spatial scale of RS imagery also plays a decisive role in FPGA design, as it makes naive full-image processing impractical. Consequently, most implementations rely on patch-based strategies, which maintain throughput but introduce an additional pre-processing overhead—often negligible compared to models’ complexity. In parallel, differences in sensing modalities also shape algorithmic design choices. For instance, HSI data, with hundreds of spectral bands, are frequently processed pixel-by-pixel (Gyaneshwar and Nidamanuri, 2022; Martins et al., 2024; Chellaswamy et al., 2024; Boyle et al., 2023; Shibi and Gayathri, 2021), while others sample channels to decrease spectral redundancies and computational load (He et al., 2023). In contrast, all studies handle SAR imagery as amplitude-only—effectively grayscale images—and therefore avoid any complex-valued operations or datatypes in the implementation.